OpenStudy (anonymous):

A certain radioactive element has a half-life of one hour. If you start with 1.0 gram of the element at noon, how much of the radioactive element will be left at 2 p.m.? 0.50 grams 0.25 grams 0.125 grams 0.0 grams

OpenStudy (chmvijay):

what you think :)

OpenStudy (anonymous):

i think its 0.25g im not to sure

OpenStudy (chmvijay):

yaaa ur right see at 12 Pm or noon 1 gram is there since 1 hr is half time where it get converted to half of initial value at 1pm = 1/2 =0.5 gram then at 2 pm 0.5/2 =0.25 gram ur right :)

OpenStudy (anonymous):

ohh hmm thought so, thank you!!(:

OpenStudy (chmvijay):

yw:)

OpenStudy (anonymous):

it is 0.25 as time is 2 hours then there are two half lives.1?2=0.5......0.5/2=0.25